### v3

# A formula for the number of up-down words

Abstract.
A word $w_{1}w_{2}\cdots w_{n}$ is said to be up-down if $w_{1}<w_{2}>w_{3}\cdots$. Carlitz and Scoville found the generating function for the number of up-down words over an alphabet of size $k$. Using properties of the Chebyshev polynomials we derive a closed-form formula for these numbers.
Keywords: up-down word, Chebyshev polynomial, distributive lattice, generating function, reflection.
Math. Subj. Class.: 68R05, 05A05, 05A15.
Department of Computer Science, Israel Academic College, 52275 Ramat Gan, Israel.
friedsela@gmail.com
1. Introduction
A permutation $\sigma=\sigma_{1}\sigma_{2}\cdots\sigma_{n}$ is said to be alternating if $\sigma_{1}>\sigma_{2}<\sigma_{3}\cdots$. By a classical result of André [1], if $c_{n}$ stands for the number of alternating permutations of length $n$, then
$\sum_{n\geq 0}c_{n}\frac{x^{n}}{n!}=\sec{x}+\tan{x}.$
The analogue question regarding words seems to have been studied first by Carlitz and Scoville [2]. Let $k\geq 1$ and $n\geq 0$ be two integers and set $[k]=\{1,2,\ldots,k\}$. A word over $k$ of length $n$ is an element of the set $[k]^{n}$. A word $w_{1}w_{2}\cdots w_{n}\in[k]^{n}$ is said to be up-down if $w_{1}<w_{2}>w_{3}\cdots$. Let $f_{k,n}$ stand for the number of up-down words over $k$ of length $n$. In [2] the generating functions for the number of up-down words over $k$ of even and odd length were found. Gao et al. [3] proved via a bijection that the number of up-down words over $k$ of length $n$ is equal to the number of order ideals of certain posets. The enumeration problem of these order ideals was addressed by Berman and Köhler [4, Example 2.3], who obtained a recursive formula for them. An evidently equivalent formulation of up-down words is that of a light ray entering a stack of glass plates such that each glass plate may reflect or transmit light. These were studied by [5].
Despite these three different formulations of the same enumeration problem, it seems that the possibility of obtaining a closed formula was overlooked. The purpose of this work is to fill this gap. We show that
$f_{k,n}=1_{n=1}+\frac{4}{2k-1}\left|\sum_{i=1}^{k-1}\frac{\sin^{2}\left(\frac{2i\pi}{2k-1}\right)}{\left(2\cos\left(\frac{2i\pi}{2k-1}\right)\right)^{n+1}}\right|,$
where $1_{n=1}$ equals $1$ if $n=1$ and 00 otherwise.
2. Main results
Let
$F_{k}(x)=\sum_{n\geq 0}f_{k,2n+1}x^{2n+1}\;\;\;\textnormal{ and }\;\;\;G_{k}(x)=\sum_{n\geq 0}f_{k,2n}x^{2n}$
be the generating functions for the number of up-down words over $k$ of odd and even length, respectively. In [2] it was shown that
$F_{k}(x)=\frac{P_{k}(x)}{Q_{k}(x)}\;\;\;\textnormal{ and }\;\;\;G_{k}(x)=\frac{1}{Q_{k}(x)},$
where
$P_{k}(x)$ $=\sum_{i=0}^{k}(-1)^{i}\binom{k+i}{2i+1}x^{2i+1},$
$Q_{k}(x)$ $=\sum_{i=0}^{k}(-1)^{i}\binom{k+i-1}{2i}x^{2i}.$
The key insight in translating these generating functions into a closed formula is their expressibility in terms of the Chebyshev polynomials. These polynomials are related to the cosine and sine functions and find use in approximation theory (e.g., [6]). They also emerge naturally in combinatorics (e.g., [7]). The Chebyshev polynomials of the first kind, denoted by $T_{n}(x)$, are defined by $T_{n}(\cos\theta)=\cos(n\theta)$ and the Chebyshev polynomials of the second kind, denoted by $U_{n}(x)$, are defined by $U_{n}(\cos\theta)\sin\theta=\sin((n+1)\theta)$. Finally, the Chebyshev polynomials of the third kind, denoted by $V_{n}(x)$, are defined by $V_{n}(\cos\theta)\cos\left(\frac{1}{2}\theta\right)=\cos\left(\left(n+\frac{1}{2}\right)\theta\right)$. With the help of these polynomials we prove the following result.
Theorem 1.
The number of up-down words over $k$ of length $n$ is given by
$f_{k,n}=1_{n=1}+\frac{4}{2k-1}\left|\sum_{i=1}^{k-1}\frac{\sin^{2}\left(\frac{2i\pi}{2k-1}\right)}{\left(2\cos\left(\frac{2i\pi}{2k-1}\right)\right)^{n+1}}\right|.$ (1)
where $1_{n=1}$ equals $1$ if $n=1$ and 00 otherwise.
Proof.
It is well known that
$T_{m}(x)$ $=m\sum_{i=0}^{m}(-2)^{i}\frac{(m+i-1)!}{(m-i)!(2i)!}(1-x)^{i},$
$U_{m}(x)$ $=\sum_{i=0}^{m}(-2)^{i}\binom{m+i+1}{2i+1}(1-x)^{i}.$
It immediately follows that
$P_{k}(x)=xU_{k-1}\left(1-\frac{x^{2}}{2}\right).$
We claim that
$Q_{k}(x)=V_{k-1}\left(1-\frac{x^{2}}{2}\right).$
Indeed, using
$T_{m}(x)$ $=\frac{1}{2}(U_{m}(x)-U_{m-2}(x)),$
$\frac{d}{dx}T_{m}(x)$ $=mU_{m-1}(x),$
$U_{m}(x)$ $=2xU_{m-1}(x)-U_{m-2}(x),$
$V_{m}(x)$ $=U_{m}(x)-U_{m-1}(x),$
we have
$Q_{k}(x)$ $=\sum_{i=0}^{k-1}(-1)^{i}\binom{k+i-1}{2i}x^{2i}$
$=k\sum_{i=0}^{k-1}(-2)^{i}\frac{(k+i-1)!}{(k-i)!(2i)!}\left(1-\left(1-\frac{x^{2}}{2}\right)\right)^{i}$
$-\sum_{i=0}^{k-1}(-2)^{i}i\frac{(k+i-1)!}{(k-i)!(2i)!}\left(1-\left(1-\frac{x^{2}}{2}\right)\right)^{i}$
$=T_{k}\left(1-\frac{x^{2}}{2}\right)-\frac{x}{2k}\frac{d}{dx}T_{k}\left(1-\frac{x^{2}}{2}\right)$
$=U_{k-1}\left(1-\frac{x^{2}}{2}\right)-U_{k-2}\left(1-\frac{x^{2}}{2}\right)$
$=V_{k-1}\left(1-\frac{x^{2}}{2}\right).$
Thus,
$F_{k}(x)=x+\frac{xU_{k-2}\left(1-\frac{x^{2}}{2}\right)}{V_{k-1}\left(1-\frac{x^{2}}{2}\right)}\;\;\;\textnormal{ and }\;\;\;G_{k}(x)=\frac{1}{V_{k-1}\left(1-\frac{x^{2}}{2}\right)}.$
We now proceed as was done in [7]. It is well-known that $V_{m}(x)$ is a polynomial of degree $m$ with $m$ distinct roots given by
$\cos\left(\frac{\left(i-\frac{1}{2}\right)\pi}{m+\frac{1}{2}}\right),\;\;i=1,\ldots,m.$
Furthermore, the coefficient of $x^{m}$ in $V_{m}(x)$ is $2^{m}$. Hence,
$V_{k-1}\left(1-\frac{x^{2}}{2}\right)$ $=2^{k-1}\prod_{i=1}^{k-1}\left(1-\frac{x^{2}}{2}-\cos\left(\frac{\left(i-\frac{1}{2}\right)\pi}{k-\frac{1}{2}}\right)\right)$
$=\prod_{i=1}^{k-1}\left(2\sin\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)-x\right)\left(2\sin\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)+x\right).$
Now, by partial fraction decomposition (e.g., [8, (5.8)]), if $q(x)$ is a polynomial of degree $m$ whose roots are all real and distinct, which factors as $q(x)=(x-\beta_{1})\cdots(x-\beta_{m})$, and $p(x)$ is a polynomial whose degree is strictly less than $m$, then
$\frac{p(x)}{q(x)}=\sum_{i=1}^{m}\frac{p(\beta_{i})}{q^{\prime}(\beta_{i})}\frac{1}{x-\beta_{i}}.$
Since
$\frac{d}{dx}U_{m}(x)=\frac{(m+1)T_{m+1}(x)-xU_{m}(x)}{x^{2}-1},$
we have
$\frac{d}{dx}V_{k-1}\left(1-\frac{x^{2}}{2}\right)=\frac{2x\left(kU_{k-2}\left(1-\frac{x^{2}}{2}\right)+(k-1)U_{k-1}\left(1-\frac{x^{2}}{2}\right)\right)}{x^{2}-4}.$
Thus,
$\frac{d}{dx}V_{k-1}\left(1-\frac{x^{2}}{2}\right)_{|x=\pm 2\sin\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)}=\pm\frac{(-1)^{i}(2k-1)}{2\cos^{2}\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)}.$
Furthermore,
$xU_{k-2}\left(1-\frac{x^{2}}{2}\right)_{|x=\pm 2\sin\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)}=\pm(-1)^{i+1}.$
We conclude that
$f_{k,n}=\frac{1}{2^{n-1}(2k-1)}\sum_{i=1}^{k-1}(-1)^{i+1}\frac{\cos^{2}\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)}{\sin^{n+1}\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)},$
if $n$ is even, and
$f_{k,n}=1_{n=1}+\frac{1}{2^{n-1}(2k-1)}\sum_{i=1}^{k-1}\frac{\cos^{2}\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)}{\sin^{n+1}\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)},$
if $n$ is odd. It follows that, for arbitrary $n$,
$f_{k,n}=1_{n=1}+\frac{4}{2k-1}\sum_{i=1}^{k-1}\frac{\cos^{2}\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)}{\left(2(-1)^{i+1}\sin\left(\frac{\left(i-\frac{1}{2}\right)\pi}{2k-1}\right)\right)^{n+1}},$
from which formula (1) easily follows. ∎
Example 2.
For $k=2$, formula (1) reduces to
$f_{2,n}=1_{n=1}+\frac{4}{3}\left|\frac{\sin^{2}\left(\frac{2\pi}{3}\right)}{\left(2\cos\left(\frac{2\pi}{3}\right)\right)^{n+1}}\right|=1_{n=1}+\frac{1}{3\cdot 2^{n-1}}\left|\frac{\left(\frac{\sqrt{3}}{2}\right)^{2}}{\left(-\frac{1}{2}\right)^{n+1}}\right|=1_{n=1}+1,$
as it should, since, in this cases, there is only one up-down word for each $n\neq 1$, namely, $121\cdots$, and, for $n=1$, there are two such words, namely $1$ and $2$.
For $k=3$, it was shown in [3, Theorem 2.3] that $f_{3,n}=F_{n+2}$, for $n\geq 2$, where $F_{m}$ stands for the $m$th Fibonacci number. This, of course, follows also from formula (1). Indeed,
$f_{3,n}$ $=\frac{4}{5}\left|\frac{\sin^{2}\left(\frac{2\pi}{5}\right)}{\left(2\cos^{n+1}\left(\frac{2\pi}{5}\right)\right)^{n+1}}+\frac{\sin^{2}\left(\frac{4\pi}{5}\right)}{\left(2\cos^{n+1}\left(\frac{4\pi}{5}\right)\right)^{n+1}}\right|$
$=\frac{1}{2}\left(\left(1+\frac{1}{\sqrt{5}}\right)\left(\frac{1+\sqrt{5}}{2}\right)^{n+1}+\left(1-\frac{1}{\sqrt{5}}\right)\left(\frac{1-\sqrt{5}}{2}\right)^{n+1}\right)$
$=\frac{1}{2}(F_{n+1}+L_{n+1})$
$=F_{n+2},$
where $L_{m}$ stands for the $m$th Lucas number (e.g., [9, pp. 132-133]).
Remark 3.
Replacing $<$ and $>$ with $\leq$ and $\geq$, respectively, in the definition of up-down words, gives rise to weakly up-down words. These seem not to have been considered at all, but turn out to be essentially the same as up-down words, since, for $n\neq 1$, the number of weakly up-down words over $k$ of length $n$ is equal to the number of up-down words over $k+1$ of length $n$. Indeed, let $w_{1}\cdots w_{n}\in[k]^{n}$ be a weakly up-down. Define an up-down word $u_{1}\cdots u_{n}\in[k+1]^{n}$ as follows:
$u_{i}=\begin{cases}w_{i},&\text{if $i$ is odd or }w_{i-1}<w_{i}>w_{i+1};\\ k+1,&\text{otherwise}.\end{cases}$
Conversely, let $u_{1}\cdots u_{n}\in[k+1]^{n}$ be an up-down word and set $u_{0}=u_{n+1}=0$. Define a weakly up-down word $w_{1}\cdots w_{n}\in[k]^{n}$ as follows:
$w_{i}=\begin{cases}u_{i},&\text{if }u_{i}<k+1;\\ \max\left\{u_{i-1},u_{i+1}\right\},&\text{otherwise}.\end{cases}$
It is easy to see that the two maps are inverse to one another and hence bijections. Thus, the number of weakly up-down words over $k$ of length $n$ is given by
$\frac{4}{2k+1}\left|\sum_{i=1}^{k}\frac{\sin^{2}\left(\frac{2i\pi}{2k+1}\right)}{\left(2\cos\left(\frac{2i\pi}{2k+1}\right)\right)^{n+1}}\right|.$
We now consider the cyclic analogue of up-down words. A word $w_{1}\cdots w_{n}\in[k]^{n}$ of even length $n$ is said to be cyclic up-down if it is up-down and $w_{n}>w_{1}$.
Theorem 4.
Let $A(x)=\sum_{n\geq 0}a_{k,2n}x^{2n}$ be the generating function for the number of cyclic up-down words of even length. Then
$A(x)=1+\frac{x^{2}}{4-x^{2}}\left(k-1+(2k-1)\frac{U_{k-2}\left(1-\frac{x^{2}}{2}\right)}{V_{k-1}\left(1-\frac{x^{2}}{2}\right)}\right).$ (2)
In particular,
$a_{k,2n}=\sum_{i=1}^{k-1}\frac{1}{\left(2\cos\left(\frac{2i\pi}{2k-1}\right)\right)^{2n}}.$ (3)
Proof.
Let us modify $F_{m}(x)$ and replace it with
$\frac{1}{x}-x+F_{m}(x)=\frac{1}{x}+\frac{xU_{m-2}\left(1-\frac{x^{2}}{2}\right)}{V_{m-1}\left(1-\frac{x^{2}}{2}\right)}.$
In particular, $f_{m,-1}=1$ and $f_{m,1}=m-1$. We count the number of cyclic up-down words over $k+1$ according to the number of occurrences and position of $k+1$. Notice that the letter $k+1$ can be placed only at an even letter index. There are $h_{k,2n}$ cyclic up-down words without any $k+1$ and $nf_{k,2n-1}$ such words with exactly one occurrence of $k+1$. Now assume that there are at least two occurrences of $k+1$ and denote by $i$ and $j$ the leftmost and rightmost index of these $k+1$s, respectively. There are $f_{k+1,2(j-i)-1}$ possibilities for the subword $w_{i+1}\cdots w_{j-1}$ and $f_{k,2n-2(j-i)-1}$ possibilities for the (cyclic) subword $w_{j+1}\cdots w_{2n}w_{1}\cdots w_{i-1}$. It follows that
$a_{k+1,2n}$ $=a_{k,2n}+nf_{k,2n-1}+\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}f_{k+1,2(j-i)-1}f_{k,2n-2(j-i)-1}$
$=a_{k,2n}+\sum_{i=1}^{n}if_{k+1,2(n-i)-1}f_{k,2i-1}.$
Multiplying both sides of the equation by $x^{2n}$ and summing over $n\geq 1$, we obtain
$A_{k+1}(x)=A_{k}(x)+\frac{x^{2}}{2}F_{k+1}(x)\left(2x+\frac{d}{dx}\left(xF_{k}(x)\right)\right).$
It follows that
$A_{k}(x)=1+\frac{x^{2}}{2}\sum_{i=1}^{k-1}F_{i+1}(x)\left(2x+\frac{d}{dx}\left(xF_{i}(x)\right)\right).$
One may show that this expression yields (2), from which the closed formula easily follows. ∎
Example 5.
For $k=2$, formula (3) should always give $1$. And, indeed,
$a_{2,2n}=\frac{1}{\left(2\cos\left(\frac{2\pi}{3}\right)\right)^{2n}}=\frac{1}{\left(-2\cdot\frac{1}{2}\right)^{2n}}=1.$
Consider now $k=3$. We have
$a_{3,2n}$ $=\frac{1}{\left(2\cos\left(\frac{2\pi}{5}\right)\right)^{2n}}+\frac{1}{\left(2\cos\left(\frac{4\pi}{5}\right)\right)^{2n}}$
$=\frac{1}{\left(2\left(\frac{-1+\sqrt{5}}{4}\right)\right)^{2n}}+\frac{1}{\left(2\left(\frac{-1-\sqrt{5}}{4}\right)\right)^{2n}}$
$=\left(\frac{1+\sqrt{5}}{2}\right)^{2n}+\left(\frac{1-\sqrt{5}}{2}\right)^{2n}$
$=L_{2n}.$
This seems to give a new combinatorial interpretation to the bisection of the Lucas numbers (cf. sequence A005248 in [10]).